# Coherent Superconductor-Semiconductor Epitaxy for Integrated Quantum Electronics

### v2

@@ Abstract
Evaluating the (zone center) phonon spectra of pure Ge and hyperdoped Ge films using Raman backscattering spectroscopy (Figure 3F), the high p-type concentration produces a new, softened mode near 280 cm-1 [20]. A consequence of this mode is the screening of the Ge-Ge optical mode, which is now coupled to the large hole concentration and explains its significant reduction in the doped specimen. Lattice dynamics calculations (Figure S11) of pure and hyperdoped systems support this observation and reveal phonon softening in hyperdoped Ge. Evaluating the Raman spectra in Figure 3F, both doped and undoped samples yield comparable Raman signals related to the Si-Si and Si-Ge-type sublattices. For the Si-Si optical vibration originating from the thin Si layers, the low dimensionality and diffusion of Ge and Ga atoms into this region lead to a significant mode softening and a strong Si-Ge band near 400 cm-1 (Figure 3C) [21]. As both doped and control films exhibit similar Si-Si, Si-Ge, and Ge-Ge bands, we infer high Ga doping does not appreciably affect the growth process.
@@
Figure 4: A 2D GISAXS pattern of the Ga:Ge epitaxial film recorded at the critical angle (0.26) of the superconducting layer. The intensity map is on a linear scale. B The out-of-plane X-ray reflectivity measurement of the Ga:Ge film with a fit made using the REFLEX software package [22]. C Cross-section of the electron scattering length density (ESLD) of the epitaxial stack derived for the fitting model of the fringes in B. D 2D GIWAXS pattern of the Ga:Ge thin film record at the critical angle (0.125). The color map is on a log scale. The family of crystallographic planes giving rise to the intense Bragg reflections of the hyperdoped layer are identified with arrows. E Corresponding integrated GIWAXS profile (qxyz) and its subsequent structural refinement (Le Bail method). The refined unit cell parameters are displayed and an expansion of the split 220 family of Bragg peaks is shown in the inset to highlight the reduced symmetry of the clamped Ga:Ge epitaxial layer in comparison to the undoped Ge control.
.3 Observation of crystalline distortion
We next resolve the microstructure of the hyperdoped heterostructure using a combination of synchrotron-based grazing incidence X-ray scattering experiments. The 2D grazing-incidence small-angle x-ray scattering (GISAXS) pattern presented in Figure 4A was recorded at the critical angle of the Ga:Ge layers (0.26 ), which is well above the critical angle of the relatively less dense Si layers (Figure S12). Pronounced Keising fringes directed in the out-of-plane direction emerge due to the interference of scattered photons within the Ga:Ge epitaxial film and are evidence of sharp interfaces between the layers. We employ X-ray reflectometry (XRR) to derive the electron density profile and compositional grading parameters (rms roughness) of the film in the normal direction. Based on the XRR fit shown in Figure 4B, compositional grading is confirmed between the layers and diffusion is revealed to be asymmetrical with respect to the stack, with larger roughness values determined for Si grown on top of Ga:Ge, compared to Ga:Ge grown on the Si layer.
@@
.4 Conclusions
In conclusion, our results establish that MBE growth of hyperdoped Ga:Ge exhibits superconductivity arising from substitutional Ga incorporation, allowing for the creation of coherent semiconductor-superconductor interfaces. TEM and EDS imaging confirm pristine structural and compositional homogeneity. Synchrotron-based methods confirm Ga occupies substitutional sites within the Ge lattice and the interplay between Ga incorporation and epitaxial constraints (substrate clamping) leads to a subtle tetragonal ($\beta$-Sn) distortion, in line with our DFT results. Electronic structure calculations show that Ga incorporation at substitutional sites shifts the Fermi level deep into the valence band and flattens the electronic bands near the R point, likely promoting superconductivity. Our findings highlight the intrinsic nature of superconductivity in hyperdoped Ga:Ge and underscore its potential as a scalable platform for epitaxial semiconductor-superconductor heterostructures using the mature group IV semiconductor toolkit.
@@ Acknowledgments
S.S.-R. acknowledges Dan Huber for assisting with TEM sample preparation. P.J. acknowledges fruitful discussions with Ross McKenzie and Ben Powell.
Funding:
P.J.S and J.S. acknowledge funding support from the United States Air Force Office of Scientific Research award number FA9550-21-1-0338. We acknowledge computational resources at the NCI National Facility through the National Computational Merit Allocation Scheme. This work was partially supported by the Australian Research Council under the following grants: DE230100173 (J.A.S.), LP210200636 (Y.-H.C., P.J.), DE220101147 (C.V.). This research was undertaken on the X-ray Absorption Spectroscopy and the SAXS/WAXS beamlines at the Australian Synchrotron, part of ANSTO. The XRR data reported in this paper were obtained at the Central Analytical Research Facility operated by Research Infrastructure (QUT). Electron microscopy was performed at the Center for Electron Microscopy and Analysis (CEMAS) at The Ohio State University.
@@
Conceptualization: P.J., J.S., and E.D. Methodology: J.A.S., P.J.S., C.V., S.S.-R. Investigation: J.A.S., P.J.S., C.V., A.B., A.D., Y.-H.C., J.V.D., S.S.-R. Visualization: J.A.S., P.J.S. Funding acquisition: J.A.S., P.J., J.S. Project administration: J.A.S., P.J.S., P.J., J.S. Supervision: J.A.S., P.J., J.S., E.D., L.W. Writing – original draft: P.J., J.A.S., P.J.S. Writing – review and editing: J.A.S., P.J.S., C.V., P.J., J.S.
@@
Data and materials availability:
Raw data available on Zenodo: doi.org/10.5281/zenodo.14782527
Supplementary materials
Materials and Methods
Supplementary Text
Figs. S1 to S14
Tables S1 to S2
References (23-0)
Supplementary Materials for
Coherent Superconductor-Semiconductor Epitaxy for Integrated Quantum Electronics
Julian A. Steele∗†, Patrick J. Strohbeen, Carla Verdi, Ardeshir Baktash,
Alisa Danilenko, Yi-Hsun Chen, Jechiel van Dijk, Lianzhou Wang,
Eugene Demler, Salva Salmani-Rezaie, Peter Jacobson, Javad Shabani
Corresponding author. Email: julian.steele@uq.edu.au, jshabani@nyu.edu
These authors contributed equally to this work.
This PDF file includes:
Materials and Methods
Supplementary Text
Figures S1 to S14
Tables S1 to S2
Materials and Methods
Thin film growth: The materials used in this study are all grown in a custom Varian Gen II molecular beam epitaxy (MBE) system on 50.8 mm undoped Ge (001) wafers. Prior to loading into vacuum, the wafers are etched in deionized water at 90 C for 15 minutes and then immediately loaded into the vacuum chamber on indium-free mounting blocks. The wafers are sequentially outgassed inside the growth reactor at 250 C, 450 C, and finally 650 C, for 15 minutes, 15 minutes, and 5 minutes, respectively, then cooled to room temperature for growth. Pure germanium (6N) and silicon (3N) source material is deposited via Thermionics HM2 e-Guns operating at a 6 kV acceleration voltage. Gallium doping is done with a standard Knudsen effusion cell source (MBE Komponenten). Atomic fluxes are measured prior to growth with a retractable quartz crystal microbalance placed in the center of the growth path in front of the substrate. The substrate is faced away from the atomic beam path(s) during flux measurement. Reported temperatures are from a thermocouple attached close to the backside of the substrate that has been calibrated to the GaAs (001) oxide-removal temperature. Samples are grown with a Ge flux of $\sim 8-8.5\times 10^{13}~{}cm^{-2},s^{-1}$, a Ga flux of $1.4-1.9\times 10^{13}~{}cm^{-2},s^{-1}$, and a Si flux of $2.5-3\times 10^{13}~{}cm^{-2},s^{-1}$. The nominally 20 nm thick films are grown as follows, as described in Ref. [16] : (1) 10 nm of Ga-doped Ge, (2) 0.5 nm of Si, (3) 10 nm of Ga-doped Ge, (4) 1 nm Si cap. Film growth is monitored using reflection high-energy electron diffraction (RHEED) along a $\langle 110\rangle$ cut (beam path $\parallel$ to $\langle 010\rangle$). Representative RHEED images throughout the growth steps are presented in Figure S1.
Four-point transport measurements: Electrical characterization is done in an Oxford TeslatronPT He4 cryostat using a HelioxVT He3 probe insert with a base temperature of $\sim$ 270 mK and magnetic field capabilities up to 12 T. Measurements are collected using a standard Van der Pauw wiring configuration on square pieces cleaved from near the center of each wafer. On-chip contacts are made via annealed In-Sn eutectic alloy at each of the four corners which are then wired to a daughterboard using gold wire.
First-principles calculations: First-principles calculations were performed using DFT as implemented in VASP [25, 26]. Projector augmented wave (PAW) pseudopotentials including semicore 3d electrons were adopted, with a plane-wave cutoff of 550 eV. A $7\times 7\times 7$ $\Gamma$-centered k-point mesh was used to sample the Brillouin zone of the conventional 8-atom unit cell of Ge (cubic diamond structure, space group $Fd\overline{3}m$). Structural optimizations were performed using the PBEsol exchange-correlation functional [27], which yields lattice parameters in excellent agreement with the experiment (see Table S1), and a convergence threshold of 0.005 eV/Å on all forces.
For the Ga-doped Ge structure with substitutional fine-structure, one host atom in the 8-atom unit cell is replaced by a Ga atom, corresponding to a doping concentration of 12.5%. Conversely, interstitial doping is modeled by placing one Ga atom at interstitial doping sites, representing a doping concentration of 11.1%. To model the effect of substrate clamping in Ga-doped Ge, the in-plane lattice parameters were fixed to the pristine Ge values, and only the out-of-plane lattice parameter was allowed to relax. K-edge spectra were simulated using the supercell core-hole method [28] within a $2\times 2\times 2$ supercell containing 64 atoms (72 atoms in the case of interstitial doping). The formation energy of a neutral Ga dopant in the Ge crystal is calculated as $E_{\textup{Form}}=E_{\textup{d}}-E_{\textup{p}}-\sum_{i}n_{i}\mu_{i}$, where $E_{\textup{d}}$ is the total energy of the structure containing the defect, $E_{\textup{p}}$ is the total energy of the pristine structure, $n_{i}$ is the number of atoms of species $i$ being added ($n>0$) or removed ($n<0$), and $\mu_{i}$ is the chemical potential. For Ga, this is the energy per atom of a Ga crystal (base-centered orthorhombic structure).
To calculate the electronic band structure and density of states, we adopted the MBJLDA meta-GGA functional [29] based on the modified Becke-Johnson exchange functional. This functional has been shown to yield excellent results for the band structure and band gap of many semiconductors, including Ge [30, 31], which standard semilocal DFT predicts erroneously to be metallic. Spin-orbit coupling was included in all electronic structure calculations. A denser $20\times 20\times 20$ k-mesh was employed to compute the density of states. To accurately determine the Fermi-level shifts in Ga-doped Ge, a $31\times 31\times 31$ k-mesh was used. Vibrational properties were computed using finite differences [32] within supercells containing 216 atoms ($3\times 3\times 3$ supercells of the 8-atom unit cell), employing $3\times 3\times 3$ $k$-point meshes and displacement amplitudes of 0.015 Å.
Scanning transmission electron microscopy: Cross-sectional specimens for scanning transmission electron microscopy (S/TEM) analysis were prepared using a focused ion beam (FIB) system (FEI Helios NanoLab 600 DualBeam). To minimize surface damage, the samples were thinned and polished with Ga ions at 5 kV. S/TEM imaging and energy-dispersive X-ray spectroscopy (EDS) were conducted using a Thermo Fisher Scientific Themis-Z microscope. The instrument operated at an accelerating voltage of 300 kV with a semi-convergence angle of 20 mrad. HAADF-STEM images were collected using a 64-200 mrad high-angle annular dark field (HAADF) detector. High-resolution images were acquired through multiple rapid scans (2048 $\times$ 2048 pixels, 200 ns per pixel), which were combined to improve the signal-to-noise ratio. EDS analysis was performed with the Super-X EDS detector, and the elemental mappings are presented as net count images.
Raman Spectroscopy: Raman spectra were recorded from the thin film surface in a backscattering configuration, using an integrated Edinburgh confocal micro-Raman instrument (RM5). Excitation was provided by one of three solid-state laser lines (532 nm, 638 nm and 785 nm) and focused on the surface using a long working lens (Olympus, 10x, 0.4 NA). Dispersion was achieved through a 1200 g/mm diffraction grating (spectral resolution 0.2 cm-1), instrument calibration was verified through checking the position of the Si band at $\pm$520.7 cm-1, and spectra were recorded using a thermoelectric CCD detector. Changes to the local vibrational structure of the hyperdoped film are first confirmed through the evaluation of the excitation energy dependence of the Raman back-scattering signals (Figure S10). At longer excitation wavelengths (785 nm), both the control and target films exhibit comparable Raman response, while more surface-sensitive information is recovered at shorter wavelengths (532 nm) to resolve differences in the thin films (Figure 3f).
Synchrotron-based GISAXS/GIWAXS: To identify the crystalline phases and microstructure of films, synchrotron-based grazing incidence wide-angle X-ray scattering (GIWAXS) data were collected at the small angle X-ray scattering/wide angle X-ray scattering (SAXS/WAXS) beamline at the Australian Synchrotron [33]: 2D scattering patterns were recorded using a wavelength of either 1.340375 Å (9.25 keV) or 0.619924 Å (20 keV), using an energy-selective Pilatus 2M detector. The measurements made using a 20 keV beam energy were to increase the observed Q-range for the structural refinements, using an energy threshold of 15 keV to suppress the X-ray fluorescence signals arising from the excited K-edges of Ge (11.1 keV) and Ga (10.37 keV). Experiments at 9.25 keV had no such issues. Two different sample-to-detector distances were employed to record both SAXS (5.5 m) and WAXS (0.45 m) profiles, each calibrated using a silver behenate reference standard. The sample and detector were enclosed in a vacuum chamber to suppress air scatter. Scattering patterns were measured as a function of the angle of incidence, with data shown acquired with an angle of incidence near the critical angle (for a given X-ray energy) to maximize scattering intensity from the sample. All collected 2D images were azimuthally integrated using PyFAI [34] and processed using a custom Python routine. The resulting unit cell models were refined using the La Bail method implemented in Fullprof [35], a comprehensive analysis software. In line with our other characterization techniques, our GIWAXS experiments failed to detect Ga droplet formation or segregation within the film. Due to the relatively small scattering volume of the Si layers, no additional peaks are indexed to this phase either.
Stress analysis: The lattice mismatch is accommodated by an elastic strain in the hyperdoped epilayer, giving a biaxial stress of $\sigma=2\mu f\cdot(1+\nu)/(1-\nu)$, where $\mu$ is the shear modulus, $\nu$ Poisson’s ratio and $f$ the misfit parameter. Assuming the hyperdoped Ga:Ge epilayer forms a cubic unit cell when unconstrained, elastic isotropy is assumed in the relevant deformation directions and the misfit parameter is given by $f=(a_{Ga:Ge}-a_{Ge})/a_{Ge}$, where $a_{Ga:Ge}$ and $a_{Ge}$ are the lattice parameters of the relaxed epilayer and substrate, respectively. We do not have such experimental data for the relaxed hyperdoped unit cell, however, since the epilayers studied here are pseudomorphically grown (Figure 3e), the values obtained from the structural refinement can be related to the relaxed lattice parameter of the Ga:Ge layer through $a_{Ga:Ge}=a_{Ge}[1+P(\Delta d/d)\perp$], where P is the elastic parameter of the films [36]. Regardless of the growth direction, the in-plane and out-of-plane biaxial strain components are given by $\epsilon_{\parallel}=(a_{Ga:Ge}-a_{Ge})/a_{Ge}$ and $\epsilon_{\perp}=\epsilon_{\parallel}(1-1/P)$, respectively. For our thin films grown on Ge (100); $P_{100}=C_{11}/(C_{11}+2C_{12})=0.5709$. To implement our calculation, we approximate the elastic constants $C_{ij}$ of the hyperdoped film to be comparable to the pure Ge crystal [36] (as there is no relevant data for GaGe alloys). Based on our hole carrier concentrations, the Ga concentration for the sample studied in Figure 3e is $\sim$11.22and has a calculated relaxed (cubic) lattice parameter of $a_{Ga:Ge}$ = 5.6614 Å, which is 0.11% larger than the Ge crystal.
Synchrotron-based EXAFS experiments: Ga K-edge X-ray absorption spectra (XAS) were captured at the Australian Synchrotron (ANSTO in Clayton, Victoria) MEX1 beamline. A double-crystal Si (111) monochromator equipped with focusing optics was used to reduce harmonic content while producing excitation energy. An inline Ga2O3 reference (the initial peak of the first derivative occurs at 10.3687 keV) was used to calibrate the monochromator at Ga-K absorption edge with E0 set to its known value of 10.3671 eV. All acquisitions were performed on thin film materials in a fluorescent mode. Brief XAS scans of specific samples were conducted to verify the stability of the materials under at least several minutes of X-ray exposure before the full acquisitions. Samples were measured with varying energy intervals over the pre-edge (5 eV) and the XANES region (0.25), with 0.035 intervals in k-space over the EXAFS. At low-k an integration time of 1 s was employed per interval, with longer integration times weighted toward high-k portions of the spectrum, up to a maximum value of 9 in k-space (max of 2 s). Multiple scans were collected at different positions of the sample. Data processing including background subtraction, scan averaging, edge-height normalization, and rebinning was performed based on the Athena program [37]. For normalization of energy spectra and removal of background, the pre-edge range was set between -150 eV and -30 eV, while the normalization range spanned from 80 to 300 eV post-edge. The order of normalization was designated as 3. For the presented EXAFS data, the k-weight was configured as 2, and the k-range for the forward Fourier transform was defined from 0.8 to 8.5.
EXAFS fine-structure modeling: Fits was conducted in the ARTEMIS, part of the IFEFFIT software package [37]. The scattering paths used to evaluate and model the data were derived from the optimized crystal structures with different defects. Using these input structures to derive the scattering paths, the only agreeable fine structure accounting for the dominant EXAFS signal arising from an R fitting window of 1.8 to 5.1 Å emerged using the calculated Ga substituted structure. To simulate the Ga substituted Ge crystal fine structure, the input paths out to 5 Å are determined using the DFT-derived Ga substitutionally doped Ge structure and the FEFF package [37], with a distance fuzz = 0.030 Å. This generated 6 unique scattering pairs centered on the Ga substitutional site, with some other lower ranked paths omitted from the actual model to avoid over-fitting (see Table LABEL:exafs-table). A common scaling factor has been used on all path distance corrections to account for the difference between the DFT simulated unit cell volume and the experiment crystal volume at room temperature. A phase correction has been made to Figure 2d in the main text using the most intense Ga-Ge single scattering path, near 2.45 Å. The data were ultimately fit using 18 variables and 26 independent points, resulting in a reasonable R-factor of 0.015. The fitting was conducted in R-space using multiple k-weightings, with k = 1, 2 and 3. Errors of individual fit parameters were determined using ARTEMIS to take into account the correlations between parameters and known parameters without error estimates were fixed during the fit.

### v3

@@ Abstract
We observe that the heavy p-type nature produces a new, softened mode near 280 cm-1 Compaan et al. (1983) in the (zone center) phonon spectra as seen in Raman backscattering spectroscopy (Figure 3f) compared to the undoped case. Details regarding this measurement are presented in the Methods and in Figure S12. This mode screens the Ge-Ge optical mode, now coupled to the large hole concentration, leading to significant reduction in intensity. Lattice dynamics calculations (Figure S13) of pure and hyperdoped systems support this observation and reveal phonon softening in hyperdoped Ge. Evaluating the Raman spectra in Figure 3f, both doped and undoped samples yield comparable Raman signals related to the Si-Si and Si-Ge-type sublattices. For the Si-Si optical vibration originating from the thin Si layers, the low dimensionality and diffusion of Ge and Ga atoms into this region lead to a significant mode softening and a strong Si-Ge band near 400 cm-1 (Figure 3c) Fournier-Lupien et al. (2013). As both doped and control films exhibit similar Si-Si, Si-Ge, and Ge-Ge bands, we infer high Ga doping does not appreciably affect the growth process. We further compute the Eliashberg spectral function Giustino (2017) (see Figure S14) yielding the total coupling strength, $\lambda=0.41$, from which we calculate $T_{\textup{c}}=0.77$ K Allen and Dynes (1975); McMillan (1968), providing evidence of a conventional phonon-mediated BCS-type pairing mechanism. This coupling primarily originates from the optical phonon modes similar to previous observations in silicon and diamond Ekimov et al. (2004); Bustarret et al. (2006). More details of this calculation are reported in the Methods section.
Dopant-driven crystalline distortion
@@
Figure 4: X-ray scattering measurement of tetragonal crystalline distortion. a 2D GISAXS pattern of the Ga:Ge epitaxial film recorded at the critical angle (0.26) of the superconducting layer. The intensity map is on a linear scale. b The out-of-plane X-ray reflectivity measurement of the Ga:Ge film with a fit made using the REFLEX software package Vignaud and Gibaud (2019). c Cross-section of the electron scattering length density (ESLD) of the epitaxial stack derived for the fitting model of the fringes in b. d 2D GIWAXS pattern of the Ga:Ge thin film record at the critical angle (0.125). The color map is on a log scale. The family of crystallographic planes giving rise to the intense Bragg reflections of the hyperdoped layer are identified with arrows. e Corresponding integrated GIWAXS profile (qxyz) and its subsequent structural refinement (Le Bail method). The refined unit cell parameters are displayed and an expansion of the split 220 family of Bragg peaks is shown in the inset to highlight the reduced symmetry of the clamped Ga:Ge epitaxial layer in comparison to the undoped Ge control.
To resolve the microstructure of the hyperdoped heterostructure we use a combination of synchrotron-based grazing incidence X-ray scattering experiments. The 2D grazing-incidence small-angle x-ray scattering (GISAXS) pattern presented in Figure 4a is recorded at the critical angle of the Ga:Ge layers (0.26), well above the critical angle of the less dense Si layers (Figure S15). Kiessig fringes due to the interference of scattered photons within the Ga:Ge film provide evidence of sharp interfaces between the layers. X-ray reflectometry (XRR) analysis is used to derive the electron density profile and compositional grading parameters (rms roughness) of the film in the normal direction. Based on the XRR fit shown in Figure 4b, compositional grading is confirmed between the layers and diffusion is revealed to be asymmetrical with respect to the stack, with larger roughness values determined for Si grown on top of Ga:Ge, compared to Ga:Ge grown on the Si layer. From the XRR fitting, we observe finite diffusion of Ge into the barrier layer, reducing the epitaxial mismatch strain and helping to preserve epitaxial coherence across this interface. Comparing the XRR electron density profile to the characteristic superconducting length scales as calculated (details in the Supplemental information) from the magnetic field dependance of this sample, presented in Figure S16, we note that the width of the diffusion profile falls well below one coherence length, $\xi$. This is in contrast to the secondary ion-mass spectroscopy (SIMS) profiles in Figure S17, which suggests a long tail of Ga diffused into the Ge substrate. We believe this disagreement to be the result of the sputtering process smearing out the backside interface during our SIMS measurements. Regardless, in the worst case scenario the Ga:Ge/substrate interface is approximately 1 $\xi$ away from the level at which the Ga dopants fall to a concentration below $1\times 10^{17}$ cm-3 (instrumental resolution). Thus, following Anderson’s theorem Anderson (1959) we do not expect significant perturbation of the observed superconducting state due to impurity diffusion since the characteristic length scale of our superconductor, $\xi$, is larger than the characteristic diffusion length scale observed in our experiments.
@@
Conclusions
Our results establish that superconductivity in MBE-grown Ga:Ge thin films arises from substitutional Ga incorporation rather than Ga-clustering, allowing for the creation of coherent superconductor-semiconductor interfaces. TEM and EDS imaging confirm pristine structural and compositional homogeneity. Synchrotron-based methods confirm Ga occupies substitutional sites within the Ge lattice and the interplay between Ga incorporation and epitaxial constraints (substrate clamping) leads to a subtle tetragonal ($\beta$-Sn) distortion, in line with our DFT results. Electronic structure calculations show that Ga incorporation at substitutional sites shifts the Fermi level deep into the valence band and flattens the electronic bands near the R point, likely promoting superconductivity. These results show significant promise for future device implementation of superconducting group IV materials; however, challenges yet remain. The band offsets of these superconducting materials interfaced with their semiconducting counterparts and the effect of long-range disorder (large kinetic inductance) in device applications are unable to be fully addressed in the current study. Looking ahead, it is necessary to further develop the proof-of-principle trilayer structures discussed here into full-fledged Josephson junction devices. Furthermore, an equivalent Ga floating layer forms during silicon doping Kesan et al. (1991) allowing for a natural extension of this study towards growth of superconducting silicon. This would enable the trade of relatively lossy germanium substrates for state-of-the-art low loss silicon substrates for use in quantum information devices. The findings presented here highlight the intrinsic nature of superconductivity in hyperdoped Ga:Ge thin films and underscore the potential of group IV covalent superconductors as a scalable platform for epitaxial superconductor-semiconductor heterostructures.
@@ Acknowledgments
S.S.-R. acknowledges Dan Huber for assisting with TEM sample preparation. P.J. acknowledges fruitful discussions with Ross McKenzie and Ben Powell. P.J.S and J.S. acknowledge funding support from the United States Air Force Office of Scientific Research award number FA9550-21-1-0338. We acknowledge computational resources at the Australian National Computational Infrastructure and Pawsey Supercomputing Research Centre through the National Computational Merit Allocation Scheme. This work was partially supported by the Australian Research Council under the following grants: DE230100173 (J.A.S.), LP210200636 (Y.-H.C., P.J.), DE220101147 (C.V.). This research was undertaken on the X-ray Absorption Spectroscopy and the SAXS/WAXS beamlines at the Australian Synchrotron, part of ANSTO. The XRR data reported in this paper were obtained at the Central Analytical Research Facility operated by Research Infrastructure (QUT). Electron microscopy was performed at the Center for Electron Microscopy and Analysis (CEMAS) at The Ohio State University.
@@
Conceptualization: P.J., J.S., and E.D. Methodology: J.A.S., P.J.S., C.V., S.S.-R. Investigation: J.A.S., P.J.S., C.V., A.B., A.D., Y.-H.C., J.V.D., F.K., A.L., D.P., S.S.-R. Visualization: J.A.S., P.J.S. Funding acquisition: J.A.S., P.J., J.S. Project administration: J.A.S., P.J.S., P.J., J.S. Supervision: J.A.S., P.J., J.S., E.D., L.W. Writing – original draft: P.J., J.A.S., P.J.S. Writing – review and editing: J.A.S., P.J.S., C.V., P.J., J.S.
@@
Methods